OpenStudy (anonymous):

find an equation of the curve that satisfies dy/dx=68yx^16 and whose y-intercept is y=5.

OpenStudy (anonymous):

1/68y dy= x^16 dx integrate

OpenStudy (anonymous):

\[1/68 \ln(y) =1/17x +C\] ............ln(y)=68/17x^17 +C

OpenStudy (anonymous):

y=e^68/17 x^17 * e^cand sub in (0,5) for x and y then solve for C